# Cristiá, Capozucca and Rossi

### v1

@@ 5.3 What verification conditions can be discharged?
  • SET is the Boolean algebra of sets. That is, a theory including hereditarily finite sets, equality, union, intersection, difference, set membership and their negations (Dovier et al.,, 2000).
  • CARD is SET extended with cardinality and combined with LIA (Cristiá and Rossi,, 2023). That is, variables denoting set cardinality can only participate in linear integer constraints. CARD cannot be combined with RIS or RA.
  • FII is CARD extended with finite integer intervals (Cristiá and Rossi, 2024c, ). As with CARD interval limits can only participate in linear integer constraints.
  • RIS is SET extended with RIS—Section 2.2. Considering RIS of the form ris(X in A,$\phi$), the theory is decidable as long as $\phi$ belongs to a decidable theory supported by $\{log\}$. When more complex RIS are used, the condition for decidability is more involved (Cristiá and Rossi, 2021b, ).
  • RA (relation algebra) is SET extended with composition, converse and the identity relation. The condition for decidability is rather involved (Cristiá and Rossi,, 2020); we give a simplified version. RA is decidable for the class of formulas not containing constraints of the form comp(R,S,R) or comp(S,R,R). As many relational constrains are defined in terms of comp (e.g., relational image), a formula not containing comp may still belong to the undecidable fragment—for example, some formulas containing (simultaneously) domain and range fall outside of the decidable fragment (Cantone and Longo,, 2014).
@@
The theory of restricted quantifiers, noted RQ, is not depicted in Figure 1. RQ is parametric w.r.t. some quantifier-free, decidable theory (Cristiá and Rossi, 2024d, )—for instance RQ could be configured with CARD or LIA or even with the decidable fragment of RA. However, the parameter theory must verify the following condition: if its solver, as a result of solving a formula, returns a conjunction of atoms including set variables, this conjunction must be satisfiable by substituting set variables by the empty set. This is not the case of CARD and FII. The decidability of RQ is as follows. Let $\{\dots/A\}$ be the domain of a RQ where $A$ is a variable. We say that $A$ is the domain variable of the RQ [^12] Clearly, if the domain of a RQ is just a variable then this is the domain variable.. The innermost formula of a RQ is called its quantifier-free formula. For instance, X is V + 1 is the quantifier-free formula of the nested RQ given in (2.4). The decidable fragment of RQ is characterized by formulas including foreach and exists with a quantifier-free formula belonging to a decidable fragment and meeting at least one of the following conditions.
@@
    The formula contains foreach and exists but does not belong to the above class (i.e., some exists occur after some foreach). In this case the condition for decidability is as follows: (i) no domain variable of a foreach is used in the quantifier-free formula; and (ii) no exists occurring after a foreach share the same domain variable. For instance, the following formula do not verify 4(ii): foreach(X in {W / A}, exists(Y in {V / A}, $\phi$)).
@@ 6 Test case generation
As we have said, the state machines specified in $\{log\}$ are considered functional prototypes. These specifications will likely be implemented in some other programming language. These implementations should be tested. Model-based testing (MBT) (Utting and Legeard,, 2006) indicates that implementations can be tested with test cases generated from their specifications. $\{log\}$ includes an implementation of the Test Template Framework (TTF) (Stocks and Carrington,, 1996), a MBT method originally developed for the Z notation. The TTF was first implemented in Java for Z specifications (Cristiá et al.,, 2014) and later connected to $\{log\}$ (Cristiá et al.,, 2013). In that implementation users are required to write a Z specification part of which is translated into $\{log\}$ and the results computed by $\{log\}$ are translated back to Z. Now the TTF has been implemented as a $\{log\}$ component that generates test cases from state machines specified in $\{log\}$. As a result we have a much simpler and cleaner tool seamlessly integrated with $\{log\}$.
@@
We will use the birthday book specification to show how the TTF works in $\{log\}$. Before applying the TTF on a specification, the specification has to be typechecked and the VCG has to be called on that specification. The TTF uses information about the state machine collected by the typechecker and the VCG. Test cases are generated for each operation in a state machine. In this example we will generate test cases for addBirthday. Test case generation proceeds by applying so-called testing tactics to the selected operation. What tactics are applied depends on the logical structure and mathematical elements used in the operation, as well as on the testing goals determined by the testing team. In $\{log\}$ the first tactic to be applied is called disjunctive normal form (DNF). The command is as follows:
[…]
addBirthday_vis    addBirthday_dnf_1    addBirthday_dnf_2addBirthday_vis is the root of the tree while addBirthday_dnf_1 and addBirthday_dnf_2 are its children—indentation is used to depict tree levels. Nodes in the testing tree are called test specifications. Each test specification is a conjunction of $\{log\}$ constraints depending on the before-state and input variables of the operation. Although we have not passed in state variables to applydnf, it uses the structure of the specification to get them. This tactic writes the precondition of the operation into DNF and generates a test specification for each term of the DNF. The following are the test specifications generated so far:
It would be possible to generate test cases from this testing tree but we will apply one more tactic in order to show how $\{log\}$ is used by the TTF more thoroughly. After DNF, tactics are applied to one or more test specifications of the testing tree. Then, we apply a second tactic known as standard partitions (SP) to the addBirthday_dnf_1 test specification. The command is as follows:
@@
{log}=> applysp(addBirthday_dnf_1,un(Known,{Name},Known_)).The first argument is the test specification and the second one is a constraint present in the operation—see addBirthdayOk in Section 3. SP defines partitions for key mathematical operators such as union, intersection, overriding, etc. Figure 2 shows the standard partition for union, intersection and set difference. The partition indicates conjunctions of conditions depending on the arguments of the operator. The rationale behind SP is that set and relational operators have non trivial implementations that deserve a thorough testing. The SP tactic generates a test specification for each conjunction defined in the partition. Then, the above command generates the following testing tree.
addBirthday_vis    addBirthday_dnf_1        addBirthday_sp_11        addBirthday_sp_12        .................   % some test specifications not shown        addBirthday_sp_17        addBirthday_sp_18    addBirthday_dnf_2addBirthday_dnf_1 is partitioned into eight test specifications due to the eight conjunctions in the standard partition for set union. Two sample test specifications are the following:
addBirthday_sp_11(Name,Date,Known,Birthday) :-  Name nin Known & Known = {} & {Name} = {}.addBirthday_sp_14(Name,Date,Known,Birthday) :-  Name nin Known & Known neq {} & {Name} neq {} & disj(Known,{Name}).Note that each test specification includes the predicate of addBirthday_dnf_1. Besides, note that the arguments of the partition ($S$ and $T$) are substituted by the arguments of the constraint passed in to the command ({Name} and Known).
@@
Finally, observe that addBirthday_sp_11 is unsatisfiable. This is a usual situation in the TTF. Given that it is impossible to generate test cases from unsatisfiable test specifications, they must be pruned from the testing tree. Detecting unsatisfiable set formulas is one of the main $\{log\}$’s capabilities. Hence, the TTF provides a command, namely prunett, that iterates over all the leaves of a testing tree and asks $\{log\}$ whether or not they are unsatisfiable. If $\{log\}$ finds an unsatisfiable test specification, prunett eliminates it from the testing tree. The resulting testing tree after calling prunett is the following:
@@
We could apply more testing tactics but that would not add anything substantial to the understanding of the $\{log\}$ framework. The final step in the TTF is generating one test case for each leaf in the testing tree. In this context a test case is an assignment of values to the before-state and input variables appearing in the test specification. In other words, a test case is a solution to some test specification. Again, $\{log\}$ is good at finding solutions to set formulas. Then, the TTF provides a command, namely gentc, that iterates over all the leaves of a testing tree and asks $\{log\}$ for a ground solution to each of them. The test cases generated by gentc are the following:
addBirthday_tc_12(Name,Date,Known,Birthday) :-  Name = name:n0 & Known = {}.addBirthday_tc_14(Name,Date,Known,Birthday) :-  Name = name:n0 & Known = {name:n1}.addBirthday_tc_2(Name,Date,Known,Birthday) :-  Name = name:n0 & Known = {name:n0}.Observe that Birthday has not been bounded to some value. This is so because no test specification includes a condition on that variable. Concerning these test cases, Birthday can assume any value. Applying SP on the postcondition un(Birthday,{[Name,Date]},Birthday_) as a third tactic would generate test cases with different values for Birthday combined with the values of these test cases. B presents the application of this tactic and the test cases so generated.
In summary, the TTF component uses $\{log\}$ as a back-end solver for automating two crucial but annoying, error-prone tasks: pruning test specifications and test case generation. Users are left with the more creative task of deciding what tactics will generate the best test suite. The current implementation of the TTF provides a total of six testing tactics; SP supports nine set and relational operators but can be easily extended by users—see $\{log\}$ user’s manual for more details (Rossi and Cristiá,, 2025, Section 13.7).
@@ 7 Concluding remarks
More advanced features should also be analyzed. For instance, first-order automated reasoning, such as the one performed by the E-prover (Schulz et al.,, 2019), may help in finding missing hypotheses. That is, $\{log\}$ could rest in E-prover to select a smaller set of promising hypotheses and then try to discharge the VC using its own engine. An even more challenging issue is to address the proofs of VC falling outside of the decision procedures available in $\{log\}$. An obvious solution would be to rest on an external interactive theorem prover (ITP). But using an external ITP brings in the problem of translating proof terms back and forth, thus introducing new sources of errors. On the other hand, even such proofs will have subgoals fitting in the decidable fragments where $\{log\}$ can be handy. Therefore, we have developed a simple prototype ITP on top of $\{log\}$ where users can call it whenever they have a subgoal fitting in some decidable fragment (Cristiá et al.,, 2022). Nevertheless, this idea needs considerably more work.

### v2

@@ 5.3 What verification conditions can be discharged?
  • SET is the Boolean algebra of sets. That is, a theory including hereditarily finite sets, equality, union, intersection, difference, set membership and their negations (Dovier et al., 2000).
  • CARD is SET extended with cardinality and combined with LIA (Cristiá and Rossi, 2023). That is, variables denoting set cardinality can only participate in linear integer constraints. CARD cannot be combined with RIS or RA.
  • FII is CARD extended with finite integer intervals (Cristiá and Rossi, 2024c ). As with CARD interval limits can only participate in linear integer constraints.
[…]
  • RA (relation algebra) is SET extended with composition, converse and the identity relation. The condition for decidability is rather involved (Cristiá and Rossi, 2020); we give a simplified version. RA is decidable for the class of formulas not containing constraints of the form comp(R,S,R) or comp(S,R,R). As many relational constraints are defined in terms of comp (e.g., relational image), a formula not containing comp may still belong to the undecidable fragment—for example, some formulas containing (simultaneously) domain and range fall outside of the decidable fragment (Cantone and Longo, 2014).
@@
The theory of restricted quantifiers, noted RQ, is not depicted in Figure 1. RQ is parametric w.r.t. some quantifier-free, decidable theory (Cristiá and Rossi, 2024d )—for instance RQ could be configured with CARD or LIA or even with the decidable fragment of RA. However, the parameter theory must verify the following condition: if its solver, as a result of solving a formula, returns a conjunction of atoms including set variables, this conjunction must be satisfiable by substituting set variables by the empty set. This is not the case of CARD and FII. The decidability of RQ is as follows. Let $\{\dots/A\}$ be the domain of a RQ where $A$ is a variable. We say that $A$ is the domain variable of the RQ [^14] Clearly, if the domain of a RQ is just a variable then this is the domain variable.. The innermost formula of a RQ is called its quantifier-free formula. For instance, X is V + 1 is the quantifier-free formula of the nested RQ given in (2.4). The decidable fragment of RQ is characterized by formulas including foreach and exists with a quantifier-free formula belonging to a decidable fragment and meeting at least one of the following conditions.
@@
    The formula contains foreach and exists but does not belong to the above class (i.e., some exists occur after some foreach). In this case the condition for decidability is as follows: (i) no domain variable of a foreach is used in the quantifier-free formula; and (ii) no exists occurring after a foreach share the same domain variable. For instance, the following formula does not verify 4(ii): foreach(X in {W / A}, exists(Y in {V / A}, $\phi$)).
@@ 6 Test case generation
As we have said, the state machines specified in $\{log\}$ are considered functional prototypes. These specifications will likely be implemented in some other programming language. These implementations should be tested. Model-based testing (MBT) (Utting and Legeard, 2006) indicates that implementations can be tested with test cases generated from their specifications. $\{log\}$ includes an implementation of the Test Template Framework (TTF) (Stocks and Carrington, 1996), a MBT method originally developed for the Z notation. The TTF was first implemented in Java for Z specifications (Cristiá et al., 2014) and later connected to $\{log\}$ (Cristiá et al., 2013). In that implementation users are required to write a Z specification part of which is translated into $\{log\}$, and the results computed by $\{log\}$ are translated back to Z. Now the TTF has been implemented as a $\{log\}$ component that generates test cases from state machines specified in $\{log\}$. As a result we have a much simpler and cleaner tool seamlessly integrated with $\{log\}$.
6.1 Brief introduction to the TTF
@@
In the TTF, test cases are generated for each operation in a state machine. Test case generation proceeds by applying so-called testing tactics to the selected operation. Each tactic partitions the input space of the operation. The input space of an operation is the set given by all the possible values that its input and before-state variables can assume. What tactics are applied depends on the logical structure and mathematical elements used in the operation, as well as on the testing goals determined by the testing team. The current implementation of the TTF provides a total of six testing tactics. Here we will discuss only two of them, namely disjunctive normal form (DNF) and standard partitions (SP). See $\{log\}$ user’s manual for more details (Rossi and Cristiá, 2025, Section 13.7).
The DNF tactic requires writing the precondition of the operation into DNF and partitioning the input space in as many sets as terms are in the DNF. In this way, the main logical alternatives of the operation are considered. For example, when DNF is applied to addBirthday it will partition the input space into a set where the new person is added to the book and another set where the person already exists and so the book remains unchanged. Hence, there will be at least one test case testing the addition of a person and another test case testing an erroneous situation. In $\{log\}$ DNF must be the first tactic to be applied.
@@
SP defines partitions for key mathematical operators such as union, intersection, overriding, etc. Figure 2 shows the standard partition for union, intersection and set difference. The partition indicates conjunctions of conditions depending on the arguments of the operator. The rationale behind SP is that set and relational operators have non trivial implementations that deserve a thorough testing. Currently, SP supports nine set and relational operators but can be easily extended by users.
@@
As a result, DNF aims at the logical structure of the operation whereas SP aims at its mathematical elements. In $\{log\}$ each testing tactic is applied by means of its own user command, as we will show in Section 6.2.
The partition of the input space is represented as a testing tree. Nodes in the testing tree are called test specifications. Each test specification is a conjunction of $\{log\}$ constraints depending on the before-state and input variables of the operation. This conjunction is called characteristic predicate of the test specification. The root of the testing tree is the input space of the operation; its characteristic predicate is truetrue—meaning that the test specification is not constrained in any way. Let $C$ and $D$ be test specifications of a given testing tree, and let $\Phi_{C}$ and $\Phi_{D}$ be their characteristic predicates. If $D$ is a child node of $C$ then $\Phi_{D}=\Phi_{C}\land\phi$ with $\phi$ being a conjunction of at least one atomic predicate different from truetrue and not present in $\Phi_{C}$. In other words, the characteristic predicate of a child node includes the characteristic predicate of its father node plus the conditions added by the testing tactic from which $D$ was created.
The testing tree is a compact representation of the status of the testing process. For example, by looking at the testing tree one can have an idea of the number of test cases that can be generated, what tactics have been applied and where.
When two or more testing tactics are applied many test specifications may turn to be unsatisfiable. Given that it is impossible to generate test cases from unsatisfiable test specifications, they must be pruned from the testing tree. Detecting unsatisfiable set formulas is one of the main $\{log\}$’s capabilities. Hence, the TTF provides a command, namely prunett, that iterates over all the leaves of a testing tree and asks $\{log\}$ whether or not they are unsatisfiable. If $\{log\}$ finds an unsatisfiable test specification, prunett eliminates it from the testing tree.
In the TTF a test case is an assignment (or binding) of values to the input and before-state variables. In other words, a test case is a solution to some test specification. Again, $\{log\}$ is good at finding solutions to set formulas. Then, $\{log\}$ provides a command, namely gentc, that iterates over all the leaves of a testing tree and asks $\{log\}$ for a ground solution to each of them. Test cases are derived only from the leaves of the testing tree because these nodes include all the conditions added by each applied tactic.
6.2 Applying the TTF to the birthday book specification
We will use the birthday book specification to show how the TTF works in $\{log\}$. Before applying the TTF on a specification, the specification has to be typechecked and the VCG has to be called on that specification. The TTF uses information about the state machine collected by the typechecker and the VCG. In this example we will generate test cases for addBirthday. As DNF is the first tactic to be applied, we issue the following command:
   {log}=> applydnf(addBirthday(Name,Date)).That is, the command waits for a term of the form $\mathit{operation(i_{1},\dots,i_{n})}$ where $i_{1},\dots,i_{n}$ are the inputs of the operation. Users decide what arguments are inputs and what are not. The resulting testing tree is shown with the writett command:
   addBirthday_vis    addBirthday_dnf_1    addBirthday_dnf_2addBirthday_vis is the root of the tree while addBirthday_dnf_1 and addBirthday_dnf_2 are its children—indentation is used to depict tree levels. The nodes of the testing tree can be exported to a file with exporttt, thus obtaining the following:
It would be possible to generate test cases from this testing tree but we can also apply more tactics to one or more of its test specifications, to further partition the input space. In this case SP is applied to the addBirthday_dnf_1 test specification with the following command:
   {log}=> applysp(addBirthday_dnf_1,un(Known,{Name},Known_)).The first argument is the test specification that we want to partition and the second one is a constraint present in the operation—see addBirthdayOk in Section 3. Then, the above command generates the following testing tree.
   addBirthday_vis    addBirthday_dnf_1    addBirthday_sp_11    addBirthday_sp_12    ................. % some test specifications not shown    addBirthday_sp_17    addBirthday_sp_18    addBirthday_dnf_2addBirthday_dnf_1 is partitioned into eight test specifications due to the eight conjunctions of the standard partition for set union—see Figure 2. Two sample test specifications are the following:
   addBirthday_sp_11(Name,Date,Known,Birthday) :-    Name nin Known & Known = {} & {Name} = {}.   addBirthday_sp_14(Name,Date,Known,Birthday) :-    Name nin Known & Known neq {} & {Name} neq {} & disj(Known,{Name}).Note that the arguments of the partition ($S$ and $T$) are substituted by the arguments of the constraint passed in to applysp (i.e., {Name} and Known).
Now we call prunett to prune possible unsatisfiable test specifications. The result is the following testing tree:
@@ 6.2 Applying the TTF to the birthday book specification
Although more tactics could be applied we stop here to keep the presentation concise. The final step in the TTF is to call gentc to generate one test case for each leaf in the testing tree. The resulting test cases that can be printed with writetc:
   addBirthday_tc_12(Name,Date,Known,Birthday) :-    Name = name:n0 & Known = {}.   addBirthday_tc_14(Name,Date,Known,Birthday) :-    Name = name:n0 & Known = {name:n1}.   addBirthday_tc_2(Name,Date,Known,Birthday) :-    Name = name:n0 & Known = {name:n0}.Observe that Birthday has not been bound to some value. This is so because no test specification includes a condition on that variable. Concerning these test cases, Birthday can assume any value. Applying SP on the postcondition un(Birthday,{[Name,Date]},Birthday_) as a third tactic would generate test cases with different values for Birthday combined with the values of these test cases. B presents the application of this tactic and the test cases so generated.
In summary, the TTF component uses $\{log\}$ as a back-end solver for automating two crucial but annoying, error-prone tasks: pruning unsatisfiable test specifications and test case generation. Users are left with the more creative task of deciding what tactics will generate the best test suite. By implementing the TTF as a $\{log\}$ component users work with only one notation and tool. Besides, it avoids costly translations between $\{log\}$ and Z which also brings in more efficiency.
@@ 7 Concluding remarks
More advanced features should also be analyzed. For instance, first-order automated reasoning, such as the one performed by the E-prover (Schulz et al., 2019), may help in finding missing hypotheses. That is, $\{log\}$ could rest on E-prover to select a smaller set of promising hypotheses and then try to discharge the VC using its own engine. An even more challenging issue is to address the proofs of VC falling outside of the decision procedures available in $\{log\}$. An obvious solution would be to rest on an external interactive theorem prover (ITP). But using an external ITP brings in the problem of translating proof terms back and forth, thus introducing new sources of errors. On the other hand, even such proofs will have subgoals fitting in the decidable fragments where $\{log\}$ can be handy. Therefore, we have developed a simple prototype ITP on top of $\{log\}$ where users can call it whenever they have a subgoal fitting in some decidable fragment (Cristiá et al., 2022). Nevertheless, this idea needs considerably more work.